OpenStudy (anonymous):

A person leaves home at 8:00 a.m. and drives to a destination at a rate of 40 mph. The person returns at a rate of 25 mph and arrives at 2:30 p.m. How far was the destination? 65 miles 100 miles 130 miles

OpenStudy (anonymous):

I also don't want a direct answer I want to know how it's done.

OpenStudy (anonymous):

he drive for 6 and half hours

OpenStudy (anonymous):

65 miles

OpenStudy (anonymous):

how did you get that @yahya90

OpenStudy (anonymous):

he drive totaly 65 miles in 6and 30 mintues

OpenStudy (anonymous):

we not depend on acceleration

OpenStudy (anonymous):

He travels for six and a half hours, which is 13/2 hours. Suppose he drives out for t hours. He then drives back for 13/2-t hours. The distance out and distance back are equal. That means\[40t=25(\frac{13}{2}-t) \implies 65t=\frac{325}{2} \implies t=\frac{5}{2} \implies d=40(\frac{5}{2})=100\]

OpenStudy (anonymous):

It would seem he went a hundred miles out and a hundred miles back. If you check with the speeds given in the problem, that should take just about six and a half hours.

OpenStudy (anonymous):

thank you @AnimalAin

OpenStudy (anonymous):

No sweat. Do math every day.